Planet X has a mass of 7.4×1026 kg and is 3.3×1010 m from Planet Y which has a mass of 9.8×1026 kg. What is the force of gravity between Planet X and Planet Y in Scientific Notation?
1
Expert's answer
2020-12-15T11:33:32-0500

We can find the force of gravity between Planet X and Planet Y from the Newton's law of universal gravitation:


"F=\dfrac{GM_XM_Y}{r^2},""F=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot7.4\cdot10^{26}\ kg\cdot9.8\cdot10^{26}\ kg}{(3.3\cdot10^{10}\ m)^2}=4.44\cdot10^{22}\ N."

Answer:

"F=4.44\cdot10^{22}\ N."